OpenStudy (anonymous):

I = $20, r = 4%, t = 2 years

OpenStudy (anonymous):

ok is there a word story with this?

OpenStudy (anonymous):

i have no idea what you are trying to solve

OpenStudy (texaschic101):

Is this a simple interest problem and your solving for p (principal) ?

OpenStudy (anonymous):

yes and im confused the answer choices are $160 $250

OpenStudy (texaschic101):

I = PRT for P you can rearrange the problem : P = I/RT now we sub P = 20 R = 4%...change to decimal form = .04 T = 2 P = I/RT P = 20/(.04)(2) P = 20 / .08 P = 250 any questions ?

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (texaschic101):

No problem :)

OpenStudy (anonymous):

i have more questions about another question

OpenStudy (texaschic101):

I can try to help...can't make no promises

OpenStudy (anonymous):

I = $450, r = 6%, t = 3 years the answer choices are $2050 and $2500

OpenStudy (texaschic101):

ok...your still looking for P so we will use P = I/RT I = 450 R = 6% -- change to decimal = .06 T = 3 lets sub now P = 450/(.06)(3) P = 450/.18 P = 2500

OpenStudy (anonymous):

I = $225, r = 5%, t = 5 years the answer choces are $900 or $9000

OpenStudy (texaschic101):

P = I/RT P = 225/(.05)(5) P = 225/.25 P = 900

OpenStudy (anonymous):

I = 126, r = 3.5%, t = 9 years the answer chioces are $40 and $400

OpenStudy (texaschic101):

P = I/RT P = 126/(.035)(9) P = 126/ .315 P = 400

OpenStudy (anonymous):

can you do 4 more please?

OpenStudy (texaschic101):

give them to me

OpenStudy (anonymous):

1.I = $364, P = $1300, r = 4% the answe choices are 6 or 7 2. I = $1275, P = $5100, r = 5% the answer choices are 4 or 5 3. I = $120, P = $400, r = 6% the chices are 5 or 6 4.I = $57, P = $380, r = 3% the choices are 4 or 5

OpenStudy (texaschic101):

1. Looking for T here .... I = PRT, change to T = I/PR T = I/PR T = 364/(1300)(.04) T = 364/52 T = 7 2. T = I/PR T = 1275/(5100)(.05) T = 1275/255 T = 5 3. T = I/PR T = 120/(400)(.06) T = 120/24 T = 5 4. T = I/PR T = 57/(380)(.03) T = 57/11.4 T = 5

OpenStudy (anonymous):

i thank you so much i wish i could pay you:) i cant than you enough you saved me like 2 hours lol

OpenStudy (texaschic101):

Glad I could help....please close this question so I do not get suspended for giving answers :)

OpenStudy (anonymous):

done:)

OpenStudy (texaschic101):

:)